OpenStudy (anonymous):

.What is the area of a circle that has a diameter of 28 m? Use 22/7 for p. [

OpenStudy (anonymous):

616cmsquared

OpenStudy (anonymous):

explain how??

OpenStudy (anonymous):

radius=diameter/2 area=piradius^2 r=28/2=14 a=22/7*14*14=616